In a certain cyclotron, a proton of mass 1.67 × 10-27 kg moves in a circle of diameter 1.6 m with an angular speed of 2.0 × 106 rad/s. What is the angular momentum of the proton?
Answer #1


radius R = 1.6/2 = 0.8 m


angular speed w = 2*10^6 rad/s

mass m = 1.67*10^-27 kg

Formula:
angular momentum L = m*R^2*w

put values

L = 1.67*10^-27*0.8^2*2*10^6

L = 2.1*10^-21 kgm^2/s ANSWER